FInd the slope if it exists, of the line containing a pair of points
(9,3) and (10,-9)

Answer by checkley77(12569)

(9,3) and (10,-9)
SLOPE=(Y2-Y1)/(X2-X1)=(-9-3)/(10-9)=-12/1=-12 ANS.